OpenStudy (anonymous):

How do you solve 1/20 + 1/45 = 1/x?

OpenStudy (anonymous):

the lcd first

OpenStudy (anonymous):

umh I have no idea.

OpenStudy (anonymous):

Wouldn't you just multiply the first equation by 45 and the 2nd equation by 20?

OpenStudy (anonymous):

x=180/13

OpenStudy (anonymous):

How'd you get that answer?

OpenStudy (anonymous):

LCD: 5?

OpenStudy (anonymous):

http://www.myalgebra.com/algebra_solver.aspx It works for everything, and you might have your own opinion on it but I dont care

OpenStudy (anonymous):

The LCD is the largest number both numbers go into. In this case, it is 180. Then write both fractions under 180: \[\frac{9}{180}+\frac{4}{180}=\frac{1}{x}\] Add up the fraction on the left side: \[\frac{13}{180}=\frac{1}{x}\] Multiply both sides by 180 and then divide by 13 to isolate \(x\) \[x=\frac{180}{13}\] Let me know if there is something you don't understand.

OpenStudy (anonymous):

How did you find the LCD?

OpenStudy (anonymous):

You are guaranteed the LCD if you multiply both denominators, like 45*20. This works but means you have to simplify your fraction at the end. You can use a factor tree or, in this case, I just added 45+45+45... until I got 180 which I know divides to 20.

OpenStudy (anonymous):

That problem went along with this word problem : It takes Lauren to finish the dishes by hand in about 30 minutes. It takes her brother Josh 45 minutes to finish them if they work together how long would it take to clean the dishes? With the answer being x=180/13. How would I write that out? How many minutes would it take them?

OpenStudy (anonymous):

\[20=2^2\times 5\] \[45=3^2\times 5\] lcm =\( 2^2\times 3^2\times 5=180\) is a sure fire method

OpenStudy (anonymous):

oh if that is the question you go right to the answer. it is \[\frac{30\times 45}{30+45}\]

OpenStudy (anonymous):

it should be 30 not 20 right?

OpenStudy (anonymous):

\[\frac{30\times 45}{30+45}=\frac{1350}{75}=18\] if that was the word problem

OpenStudy (anonymous):

whoops I meant 20 minutes

OpenStudy (anonymous):

wellll in that case it is \[\frac{20\times 45}{20+45}=\frac{900}{65}=\frac{180}{13}=13\tfrac{11}{13}\]